1.A positive charge of 10 -8 C experiences a force of 0.25 N when located at a certain point in an electric field. Find the intensity of the electric field at that point.

1
Expert's answer
2022-03-20T18:48:30-0400

The intensity of the electric field

"E=\frac{F}{q}=\rm \frac{0.25\: N}{10^{-8}\: C}=2.5\times 10^7\: N/C"